OpenStudy (anonymous):

The border line of the linear inequality 3x+7y >or equal to 10 is dotted. A. True B. False

OpenStudy (anonymous):

\(>\) or \(<\) dotted,

OpenStudy (anonymous):

\(\le\) or \(\ge\) solid

OpenStudy (anonymous):

The equateion should be 3x+7y is greater than or equal to 10 is dotted. True or false?

OpenStudy (anonymous):

@satellite73 it is ≥ :)

OpenStudy (anonymous):

oh i see it changed from \(>\) to \(\ge\)

OpenStudy (anonymous):

so it should be solid and not dotted

OpenStudy (anonymous):

I didnt know how to type it so I put the sign and then "or equal to" :p

OpenStudy (anonymous):

ok i didn't know how to read so i ignored that part

OpenStudy (anonymous):

@satellite73 I posted a picture for you to click on of the question if that makes it easier :)

OpenStudy (anonymous):

as they say in perry mason, asked and answered

OpenStudy (anonymous):

\(\ge\) solid \(>\) dotted

OpenStudy (anonymous):

of course :)

OpenStudy (anonymous):

so the answer would be false? @satellite73

OpenStudy (anonymous):

false for sure

OpenStudy (anonymous):

thank you!! :) @satellite73